LBTO UT Apr 20: INAF Service and SO Training

(All times are MST)

<a name="Night_crew"></a> Night crew:

Mountain: Steve A.

Tucson: Michelle, Shane, Andrew

<a name="Summary"></a> Summary

LBTO Service Observing for INAF with LUCI2. Started the night completing last 2 scripts for medium priority Polletta MOS mask B in very good seeing (0.5"-0.7") and very light cirrus. As clouds absolutely cleared, moved to highest priority Guaita imaging program in excellent seeing 0.5-0.8 and finished 3/7 Br-gamma imaging scripts and ~1/7 H2 scripts before and standard. Finished the night with high priority Beltran program. Finished A2 target in H2 and Ks in 0.8" seeing

<a name="Nightlog"></a> Calibrations

10 of 6 x 5.0, LIR, INTEGRATED for De Barros, Polletta

10 of 12 x 10.0, LIR, INTEGRTED for Beltran

<a name="Nightlog"></a> Nightlog

18:59 Sunset.

19:03 Issue with the left side M2, Steve is looking into it.

19:42 Slewing to Polletta Mask B Telluric target.

19:45 Starting acquisition of Polletta Mask B Telluric, instrument configuring.

19:50 12 deg twilight.

19:52 Initial offset was -33.14, -1.85. 0\x94.6 on guider.

20:03 Telluric very bright and seeing is very good; not in linear range. Editing to DIT = 3s and trying again.

20:06 Reapplying offset from earlier and going directly to through slit image.

20:08 Taking Polletta Mask B Telluric

20:16 Slewing to Polletta Mask B the science field.

20:18 Starting acquisition of Polletta Mask B. 0\x94.55 on guider.

20:35 Starting science. Offset used was 3.25, -3.00, 0.0062 deg.

20:42 Shell RIP, img #46 questionable

20:47 Recovered, repeating step 46 having first reversed the offset.

20:56 0\x94.7 on guider.

21:34 Repeating script for Polletta Mask B. Note to INAF: Each 45 minute exposure time takes about 60 minutes

22:04 0\x94.6 on guider.

22:27 Repeating 5 frames for Polletta Mask B. 0\x94.7 on guider.


Summary: Polletta Mask B should be done between 3/5 sets on 20160420 UT and tonight 2/5. We realized the mistake we found in offsets on Mask C is not present in Mask B scripts. The ending point after offsets is -1 on Mask C but 0 on Mask B. This meant that our attempt to correct the error last night may have pushed a few frames close to the edge of the slits than we'd like. The extra 5 frames should make up for that.



22:41 Slewing to Guaita CD Target, BrGamma

22:48 Guide star is double, seeing is good so we\x92ll start imaging while we find a new one and edit the script.

22:51 Image quality looks excellent. We will look for an alternative guide star, but will not change the script unless we see an issue with image quality. Guider seeing estimate is not reliable given the double star, DIMM seeing is 0\x94.8. I note that the DIMM consistently reads worse than the guider. On image 0\x94.45, filter is BrGa.

23:33 Running second Guaita CD Target, BrGamma script. DIMM reading 0\x94.75, 0\x94.45 on image.

00:08 Seeing is trending up on the DIMM, but we still look ok on the images.

00:13 0\x94.65 on image, 1\x94 on DIMM. Seeing is trending up a little, but is still good.

00:21 Starting third Guaita CD Target, BrGamma sequence.

00:32 image FWHM ~ 0\x94.55

00:57 Seeing is softening a little 1\x94.2 on DIMM, 0\x94.8 on image.

01:10 Slewing to Guaita Standard P266-C

01:12 Checking pointing. Seeing trending back down, 0\x94.87 on the DIMM

01:14 Guide star is a galaxy! Slewing instead to Guaita Standard P212-C

01:19 Img #179 bad, script re-loaded too early

01:25 This standard is a bit low, Guaita Standard P172-E is higher, slewing to it, but script is incorrect. Can execute Br-gamma

01:56 Slewing back to Guaita CD Target H2. Completed ~20/30 frames.

02:02 Seeing on image ~ 0\x94.7

Summary: Guaita CD Target Br-gamma is 3/7 complete and about H2 1/7 complete. Seeing softened in the middle and we did the standard but then it improved again and we were able to go back and get some H2 data as well.

02:36 Slewing to Beltran standard S875-C, H2; INAF: this required changing the guide star to something brighter

02:54 Starting Beltran standard S875-C Ks INAF: this required changing the guide star to something brighter

03:03 Slewing to Beltran Field A2, H2. Running 15 steps to complete the set begun on 20160419.

03:55 0\x94.66 on guider.

03:58 Running Ks observations on the A2 field.

04:25 Seeing on guider is 0\x94.7

04:36 Slewing Beltran science, Field A4, Ks 0\x94.73 on guider..

04:41 0\x94.7 on first exposure.

04:47 12 deg twilight.

04:53 Img #275, taken with tracking only. The offset was too large for the guide probe to reach the required position. Image quality looks fine.

04:58 The sky is starting to brighten, we may not be able to complete this sequence. We will continue for as long as we can guide.

05:11 285 is being taken with tracking only due to the same guide probe issue. A little more astigmatism, but 0\x94.6 on exposure.

05:14 Finished on sky for the night. Steve is closing up. We will begin calibrations once he is finished.

05:28 Running Brg internal flats. About 12k ADU per flat.

05:37 Running H2 internal flats. About 12k ADU per flat.

05:38 Sunrise.

05:42 Ks internal flats. ABout 10k ADU per flat.

05:47 Running spectroscopic flats and arcs for Polletta field B. Levels look good, around 8k ADU in the flats.

06:02 Calibrations are complete. Setting LUCI2 to its end of night position

Topic revision: r2 - 21 Apr 2016, AndrewCardwell
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback